OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 10 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice. Please Use Substitution

OpenStudy (anonymous):

x - y = 10 2x + y = 2 rearranging the first equation by taking -y to the other side: x = 10 + y since we have found what x is we can substitute it in the other equation like so: 2(10 + y) + y = 2 20+2y+y = 2 20+3y = 2 3y = 2-20 3y = -18 y = -18/3 y = -6 and since we know x = 10 + y x = 10 - 6 x = 4

OpenStudy (anonymous):

And the second part, I would've used elimination however since you wanted me to use substitution, I would say that I choose substitution because by looking at the first equation I can easily rearrange the equation and plug it as there aren't very big numbers, Moreover it would consume less time than canceling out the equations.